OpenStudy (jjenk):

Let f(x) = x + 8 and g(x) = x - 4. What is f(x) • g(x)? Help please

jimthompson5910 (jim_thompson5910):

assuming that's a multiplication symbol, you just multiply the two functions

jimthompson5910 (jim_thompson5910):

f(x) * g(x) = (x+8)*(x-4) f(x) * g(x) = x*(x-4) + 8*(x-4) ... use the distributive property f(x) * g(x) = x*x-x*4 + 8*x-8*4 ... use the distributive property again f(x) * g(x) = x^2 - 4x + 8x - 32 f(x) * g(x) = x^2 + 4x - 32

jimthompson5910 (jim_thompson5910):

hopefully that makes sense

OpenStudy (anonymous):

do you still need help?